40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 14 03:42:31 2015, Jenne, Update, LSC, Thoughts on going forward with variable finesse 
    Reply  Thu Jan 15 04:41:01 2015, Jenne, Update, LSC, Thoughts on going forward with variable finesse 
       Reply  Thu Jan 15 14:28:14 2015, Jenne, Update, LSC, LSC model change idea SwitchableErrorSignalProposal_Jan2015.pdf
          Reply  Fri Jan 16 03:31:35 2015, Jenne, Update, LSC, LSC model change implemented LSCorig_15Jan2015.pngLSCpost_15Jan2015.png
             Reply  Fri Jan 16 18:46:15 2015, Koji, Update, LSC, LSC model change implemented 
                Reply  Fri Jan 16 20:01:32 2015, Jenne, Update, LSC, LSC model change implemented 
             Reply  Tue Jan 20 15:09:01 2015, Jenne, Update, LSC, LSC model change implemented 
             Reply  Sat Mar 7 18:53:48 2015, Jenne, Update, LSC, DoF selector matrices replaced with filter banks DoFselectors_7March2015.png
Message ID: 11113     Entry time: Sat Mar 7 18:53:48 2015     In reply to: 10910
Author: Jenne 
Type: Update 
Category: LSC 
Subject: DoF selector matrices replaced with filter banks 

This is work that I did yesterday but didn't have time to elog.  Since it seems non-trivial to give ourselves ramping matrices, but we only really needed the ramping in the DoF selector matrices, I've replaced the separate _A and _B parts with full filterbanks.  Recall from elog 10910 that I had given each degree of freedom's _A and _B input options an offset, an epics monitor and a test point.  Now those are removed, and handled inside of the filter banks.  The outputs of the filter banks sum together.

This required some screen modifications, but everything should work the same way that it did before this change.  I've also changed the DAQ channels from the _A_ERR and _B_ERRs that I had hand-created to now be the _A_OUT and _B_OUT test points from the filter banks (acquired at 2048Hz). 

I have not yet modified the burt snapshots for the ifo configure screen.  The arms will work the same as always, since they didn't have any selector matrix stuff ever, but the rest still need tweaking.

Attachment 1: DoFselectors_7March2015.png  301 kB  | Hide | Hide all
DoFselectors_7March2015.png
ELOG V3.1.3-